OpenStudy (anonymous):

Solve. x/3=x+2/2 x= Answer choices -6 -2 2

OpenStudy (anonymous):

\[\frac{ x }{ 3 }=\frac{ x+2 }{ 2 }\] or \[\frac{ x }{ 3 }=x+\frac{ 2 }{ 2 }\]

OpenStudy (xapproachesinfinity):

-6

OpenStudy (anonymous):

@xapproachesinfinity we are not suppose to just give the answer we are to help them and guide them so they can figure it out themselves

OpenStudy (xapproachesinfinity):

You are right! I actually did with some guys, but I'm still new adn it takes me time using the editor lol

OpenStudy (anonymous):

can you explain how you got -6 please

OpenStudy (xapproachesinfinity):

x/3=(x+2)/2 i multiply both sides by 2 i get 2/3x=x+2 i substract x from both sides i get -1/3x=2 then multiply by -3 i get x =-6

OpenStudy (anonymous):

thank you

OpenStudy (xapproachesinfinity):

you are welcome

OpenStudy (anonymous):

Thank you!!

OpenStudy (xapproachesinfinity):

pleasure!